### v1

# Effective Bertini theorems and zeros of $p$-adic forms of degree 7

Abstract.
We establish an effective Bertini-type theorem for hypersurfaces $X_{f}\colon f=0$ defined over a finite field $k$ for which $f$ has no linear factors over the algebraic closure $\overline{k}$. Given a line $L$ defined over $k$ and a nonreduced $\overline{k}$-point $x$ on $X_{f}\cap L$, we give an upper bound on the number of planes $P$ containing $L$ for which $X_{f}\cap P$ contains a line through $x$. Underlying this result is a factorization algorithm for bivariate polynomials originally due to Kaltofen, which we present with slightly relaxed hypotheses. Our primary application is to Artin’s conjecture on $p$-adic forms of degree 7: if $K/\mathbb{Q}_{p}$ is a finite extension with residue field isomorphic to $\mathbb{F}_{q}$ and $F(x_{0},\ldots,x_{n})\in K[x_{0},\ldots,x_{49}]$ is a homogeneous form of degree 7, then there exists a $K$-solution to $F=0$ whenever q>679q>{679}. This improves on a result of Wooley.
1. Introduction
In this work, we prove an effective Bertini-type theorem with a focus on ruling out certain factorizations in planar slices of a hypersurface in $\mathbb{P}^{n}$. Over a finite field $\mathbb{F}_{q}$ for $q$ sufficiently large, this ensures existence of a plane with the desired properties. We give an application of this to a conjecture of Artin concerning zeros of degree seven forms over $p$-adic fields.
1.1. On Bertini theorems
The classical Bertini theorem states that given a smooth projective variety $X\hookrightarrow\mathbb{P}^{n}$ over a field $k$, a generic hyperplane $H$ of $\mathbb{P}^{n}$ has the property that $H\cap X$ is smooth. In fact, there are several versions of this type of theorem where the adjective “smooth” is replaced with some other property of $X$, and a sufficiently general hyperplane intersected with $X$ is guaranteed to have the same property. If $k$ is infinite, the genericity of such a hyperplane is sufficient to guarantee the existence of one such hyperplane, but when $k=\mathbb{F}_{q}$ this need not be the case. One way to handle this is to replace the hyperplane $H$ by a hypersurface of higher degree: Poonen showed for $X$ a smooth quasiprojective subscheme of $\mathbb{P}^{n}$, a positive density of hypersurfaces have smooth intersections with $X$ [Poo04].
In another direction, Kaltofen [Kal95] gives an effective Bertini theorem replacing smoothness with geometric irreducibility for planar slices: given a geometrically irreducible hypersurface $X\subset\mathbb{P}^{n}$ for $n\geq 3$, he gives an upper bound on number of planes $P\subset\mathbb{P}^{n}$ for which $X\cap P$ is not geometrically irreducible in terms of the degree of $X$. Over a finite field $\mathbb{F}_{q}$, this is essentially an effective version of the Hilbert Irreducibility Theorem. Cafure and Matera [CM06] modified this approach further to bound how often such a planar slice contains irreducible components of small degree. We give a modest improvement on this result below.
Theorem A (See Theorem 2.1 and [CM06, Corollaries 3.2, 3.4]).
Suppose $f(x_{1},\ldots,x_{n})\in\mathbb{F}_{q}[x_{1},\ldots,x_{n}]$ is an absolutely irreducible (not necessarily homogeneous) polynomial of total degree $d$. Define a family of bivariate polynomials
$f_{\mathbf{v},\mathbf{w},\mathbf{z}}=f(X+v_{1},w_{2}X+z_{2}Y+v_{2},\ldots,w_{n}X+z_{n}Y+v_{n})\in\mathbb{F}_{q}[X,Y]$
for
$(\mathbf{v},\mathbf{w},\mathbf{z})=(v_{1},\ldots,v_{n},w_{2},\ldots,w_{n},z_{2},\ldots,z_{n})\in\mathbb{F}_{q}^{3n-2}.$
  1. (i)
    There exist at most
    $\frac{1}{8}\left(3d^{4}-2d^{3}+13d^{2}+2d\right)q^{3n-3}$
    tuples $(\mathbf{v},\mathbf{w},\mathbf{z})\in\mathbb{F}_{q}^{3n-2}$ such that $f_{\mathbf{v},\mathbf{w},\mathbf{z}}$ is not absolutely irreducible.
  2. (ii)
    Fix a positive integer $D<d$. Then there exist at most
    $\frac{d}{8}\left(-D^{4}+4D^{3}d-6D^{3}+12D^{2}d-11D^{2}+8Dd-6D+16d\right)q^{3n-3}$
    tuples $(\mathbf{v},\mathbf{w},\mathbf{z})\in\mathbb{F}_{q}^{3n-2}$ for which $f_{\mathbf{v},\mathbf{w},\mathbf{z}}$ has an irreducible factor of degree at most $D$ over $\overline{\mathbb{F}_{q}}$.
We modify Theorem A further to address the situation where $f$ may be reducible over $\overline{\mathbb{F}_{q}}$, but possesses no linear factors, subject to the hypothesis that for some $(\mathbf{v}_{0},\mathbf{w}_{0})\in\mathbb{F}_{q}^{2n-1}$ we have $f_{\mathbf{v}_{0},\mathbf{w}_{0},\mathbf{z}}(X,0)$ has a simple root at $\alpha_{0}\in\overline{\mathbb{F}_{q}}$. More precisely, we construct a polynomial $\Psi(Z_{2},\ldots,Z_{n})$ which, if nonzero, implies that $f_{\mathbf{v}_{0},\mathbf{w}_{0},\mathbf{z}}$ has no linear factor vanishing at $(\alpha_{0},0)$; for the full statement, see Theorem 2.6. In particular, we obtain the following corollary.
Theorem B (See Corollary 2.7).
Suppose $f(x_{1},\ldots,x_{n})$ is a (not necessarily homogeneous) polynomial of total degree $d$ over $\mathbb{F}_{q}$, possessing no linear factors over $\overline{\mathbb{F}_{q}}[x_{1},\ldots,x_{n}]$. Moreover, suppose there exist $(\mathbf{v}_{0},\mathbf{w}_{0})\in\mathbb{F}_{q}^{2n-1}$ such that $f_{\mathbf{v}_{0},\mathbf{w}_{0},\mathbf{z}}(X,0)\in\overline{\mathbb{F}_{q}}[X]$ is degree $d$ with a simple root $\alpha_{0}\in\overline{\mathbb{F}_{q}}$.
Then there exist at most $(3d-3)q^{n-2}$ tuples $\mathbf{z}\in\mathbb{F}_{q}^{n-1}$ such that $f_{\mathbf{v}_{0},\mathbf{w}_{0},\mathbf{z}}$ has no linear factor over $\overline{\mathbb{F}_{q}}$ containing $(\alpha_{0},0)$.
Our main application is to Artin’s conjecture for $p$-adic forms of degree 7.
1.2. On Artin’s conjecture
Let $\mathbb{Q}_{p}$ denote the field of $p$-adic numbers and let $K/\mathbb{Q}_{p}$ be a finite extension with residue field $\mathbb{F}_{q}$. Artin conjectured that $K$ is $C_{2}$; that is, for all positive integers $d$ and all $n\geq d^{2}$, every homogeneous polynomial $F\in K[x_{0},\ldots,x_{n}]$ of degree $d$ has a nontrivial zero in $K$. This conjecture was proven false by Terjanian, who found a counterexample with $d=4$ and $n=17$ over $K=\mathbb{Q}_{2}$ [Ter66]; infinitely many counterexamples have since been shown to exist [LM83].
However, weaker versions of the conjecture are known to hold. For a fixed degree $d$, we say that $K$ is $C_{2}(d)$ if for all $n\geq d^{2}$, every degree $d$ form $F\in K[x_{0},\ldots,x_{n}]$ has a nontrivial zero in $K$. All $p$-adic fields $K$ are $C_{2}(2)$ and $C_{2}(3)$, due to Hasse and Lewis, respectively [Lew52]. Using techniques from logic, Ax and Kochen showed that $p$-adic fields are asymptotically $C_{2}(d)$, in the sense that for a fixed degree $d$, $\mathbb{Q}_{p}$ is $C_{2}(d)$ for all but finitely many primes $p$ [AK65]. Moreover, a finite extension $K/\mathbb{Q}_{p}$ is $C_{2}(d)$ whenever $p\gg_{d,[K:\mathbb{Q}_{p}]}0$. In fact, all of the known counterexamples to Artin’s conjecture involve composite degrees $d$, divisible by q1q-1, leaving open the following conjecture.
Conjecture (Artin’s conjecture for prime degree $p$-adic forms).
Let $d>3$ be prime and $K/\mathbb{Q}_{p}$ be a finite extension. Then for all $n\geq d^{2}$, every degree $d$ form $F\in K[x_{0},\ldots x_{n}]$ has a zero in $K$, i.e. $K$ is $C_{2}(d)$.
In degree $d=7$, we improve the best known effective lower bound on $q$ for which Artin’s conjecture holds.
Theorem C.
Let $K/\mathbb{Q}_{p}$ be a $p$-adic field with residue field $\mathbb{F}_{q}$. If q>679q>{679}, then for all $n\geq 49$, every degree 7 form $F\in K[x_{0},\ldots,x_{n}]$ has a zero in $K$, i.e. $K$ is $C_{2}(7)$.
To establish Theorem C, we follow a $p$-adic minimization approach due to Birch and Lewis [BL59] and later refined by Laxton and Lewis [LL65] to give effective bounds on $q$ for which Artin’s conjecture holds in degrees $d=5,7,11$. They proposed to reduce to the case of so-called reduced forms $F$, then use an effective version of the Lang–Weil bound for the number of $\mathbb{F}_{q}$-points on a variety to produce a nonsingular zero of $\overline{F}$, the image of $F$ under the residue map, when $q$ is sufficiently large. This then lifts to a $K$-zero of $F$ by Hensel’s lemma.
Refinements of this general strategy have shown that $K$ is $C_{2}(5)$ whenever q>5q>{5} due to [LY96, HB10, Dum17] and to forthcoming work of the authors, and that $K$ is $C_{2}(7)$ (resp. $C_{2}(11)$) whenever q>883q>{883} (resp. q>8053q>{8053}) due to [Kna01, Woo08]. This strategy breaks down for prime degrees $d>11$, since $d$ can be written as the sum of composite numbers, which allows for the hypersurface given by $\overline{F}=0$ to have only totally nonreduced nonlinear components.
In degree 5, Leep and Yeomans [LY96] refined the $p$-adic minimization techniques of Birch and Lewis by exploiting the structure of singular plane quintic curves. They argue that if $\overline{F}$ has no nonsingular zeros (which would then be liftable to $K$ by Hensel’s lemma), then there exists a plane on which $\overline{F}$ restricts to a quintic curve with at least 3 singularities. The presence of these singularities reduces the $q$ threshold for which explicit point counting methods guarantee a liftable point.
In contrast, in degree 7 (resp. 11), Wooley’s bound [Woo08, Theorem 1] is obtained using the effective Bertini theorems [CM06, Corollaries 3.2, 3.4] to find plane curves of degree 7 (resp. 11) with a geometrically integral component, then applying the Hasse–Weil bounds. In particular, Wooley does not exploit any singularities in the relevant degree 7 (resp. 11) curves, citing the absence of rigidity on the anatomy of a polynomial imposed by the relatively low degree of a quintic form.
Inspired by the work of Leep and Yeomans in the quintic case, we further exploit the structure of reduced degree 7 forms either to find nonsingular zeros of $\overline{F}$, or to show that (a dehomogenization of) $\overline{F}$ satisfies the hypotheses of Theorem B. Upon applying Theorem B, we produce a plane curve with at least two singular $\mathbb{F}_{q}$-points and a geometrically irreducible component; these extra singularities lower the $q$ threshold for which we can find nonsingular points on the curve, and hence zeros of $F$ in $K$.
1.3. Notation and organization
We reserve the following to have consistent meaning throughout:
  • $p$ denotes a prime;
  • $q$ denotes a power of $p$;
  • $K$ denotes $p$-adic field with residue field $\mathbb{F}_{q}$;
  • $\overline{k}$ denotes the algebraic closure of a field $k$;
  • $F$ denotes a homogeneous form over $K$.
We will often write $f=\overline{F}$ for the image of $F$ under the residue map, to avoid confusion with the algebraic closure.
This paper is organized as follows. In Section 2 we describe a factorization algorithm for bivariate forms, which we use to prove Theorem B. We give a proof of Theorem A in Section 2.3. In Section 3 we record some facts about degree 7 forms over $\mathbb{F}_{q}$ and in Section 4 we recall the definition of reduced forms and collect auxiliary results about them. Finally, we prove Theorem C in Section 5.
Acknowledgements
This work began during a Research in Groups workshop at the International Centre for Mathematical Sciences in Edinburgh in June 2024. Further work was undertaken during Research in Pairs meetings at Banff International Research Station in December 2024 and the Centre International de Rencontres Mathématiques in Luminy in June 2025. The authors gratefully acknowledge the support of ICMS, BIRS, and CIRM for fostering a stimulating research environment, instrumental to the development of this work. The authors would also like to thank Tim Browning, Roger Heath-Brown, David Leep, Bjorn Poonen, and Trevor Wooley for thoughtful comments on an earlier draft.
LB is grateful for the support of the U.S. National Science Foundation (DMS-2418835) and the Simons Foundation (MPS-TSM–00007992). CK was supported by the Additional Funding Programme for Mathematical Sciences, delivered by EPSRC (EP/V521917/1) and the Heilbronn Institute for Mathematical Research, and an AMS-Simons Travel Grant.
2. Effective Bertini theorems
Let $k$ be an arbitrary field and suppose $f(x_{0},\ldots,x_{n})$ is a homogeneous polynomial of degree $d$ over $k$, irreducible over $\overline{k}$. Recall that generically, the intersection of the projective hypersurface $X_{f}\colon f=0$ remains geometrically irreducible when intersected with a hyperplane; this is Bertini’s theorem for irreducibility. However, when $k=\mathbb{F}_{q}$, this does not suffice to guarantee the existence of even one such hyperplane.
By an effective Bertini theorem, we refer to a result which bounds how often the non-generic behavior occurs. When $k=\mathbb{F}_{q}$, this can then be used give a lower bound on $q$ for which the generic behavior is guaranteed to occur at least once. We will be particularly interested in intersecting $X_{f}$ with a plane, rather than a hyperplane, and find it convenient to work in affine coordinates.
Dehomogenizing by setting $x_{0}=1$, consider the family of bivariate polynomials given by
$f_{\mathbf{v},\mathbf{w},\mathbf{z}}=f(1,X+v_{1},w_{2}X+z_{2}Y+v_{2},\ldots,v_{1},w_{n}X+z_{n}Y+v_{n})\in k[X,Y]$
for
$(\mathbf{v},\mathbf{w},\mathbf{z})=(v_{1},\ldots,v_{n},w_{2},\ldots,w_{n},z_{2},\ldots,z_{n})\in k^{3n-2}.$
We think of $f_{\mathbf{v},\mathbf{w},\mathbf{z}}$ as cutting out an affine plane curve, obtained by slicing the $x_{0}\neq 0$ affine patch of $X_{f}$ by the plane determined by $(\mathbf{v},\mathbf{w},\mathbf{z})$. If the original $f(x_{0},\ldots,x_{n})$ is absolutely irreducible, then generically so is $f_{\mathbf{v},\mathbf{w},\mathbf{z}}$.
This was made effective by Kaltofen, using an algorithm for testing irreducibility of a bivariate polynomial to deduce the existence of a polynomial condition on $(\mathbf{v},\mathbf{w},\mathbf{z})$ whose vanishing is implied by reducibility of $f_{\mathbf{v},\mathbf{w},\mathbf{z}}$ [Kal95]. Cafure and Matera modified this approach to detect how often $f_{\mathbf{v},\mathbf{w},\mathbf{z}}$ has factors of small degree [CM06]. We recall Theorem A, a mild strengthening of [CM06, Corollaries 3.2, 3.4]. For completeness, we describe the necessary modifications to the proofs of [CM06, Corollaries 3.2, 3.4] in §2.3.
Theorem 2.1 (See [CM06, Corollaries 3.2, 3.4]).
Suppose $f(x_{1},\ldots,x_{n})\in\mathbb{F}_{q}[x_{1},\ldots,x_{n}]$ is an absolutely irreducible (not necessarily homogeneous) polynomial of total degree $d$.
  1. (i)
    There exist at most
    $\frac{1}{8}\left(3d^{4}-2d^{3}+13d^{2}+2d\right)$
    tuples $(\mathbf{v},\mathbf{w},\mathbf{z})\in\mathbb{F}_{q}^{3n-2}$ such that $f_{\mathbf{v},\mathbf{w},\mathbf{z}}$ is not absolutely irreducible.
  2. (ii)
    Fix a positive integer $D<d$. Then there exist at most
    $\frac{d}{8}\left(-D^{4}+4D^{3}d-6D^{3}+12D^{2}d-11D^{2}+8Dd-6D+16d\right)q^{3n-3}$
    tuples $(\mathbf{v},\mathbf{w},\mathbf{z})\in\mathbb{F}_{q}^{3n-2}$ for which $f_{\mathbf{v},\mathbf{w},\mathbf{z}}$ has an irreducible factor of degree at most $D$ over $\overline{\mathbb{F}_{q}}$.
Example 2.2.
Let $d=7$ and assume $f(x_{1},\ldots,x_{n})$ is irreducible over $\overline{\mathbb{F}_{q}}$. By Theorem 2.1(i), when q>896q>{896} we have that there must exist at least one $(\mathbf{v},\mathbf{w},\mathbf{z})\in\mathbb{F}_{q}^{3n-2}$ such that $f_{\mathbf{v},\mathbf{w},\mathbf{z}}(X,Y)$ is absolutely irreducible. Suppose instead that we only require that $f_{\mathbf{v},\mathbf{w},\mathbf{z}}(X,Y)$ does not have a linear factor over $\overline{\mathbb{F}_{q}}$. Then applying Theorem 2.1(ii) with $D=1$, such a $(\mathbf{v},\mathbf{w},\mathbf{z})$ exists whenever q>224q>{224}.
The latter condition corresponds to the plane curve cut out by $f_{\mathbf{v},\mathbf{w},\mathbf{z}}$ having no linear irreducible components over $\overline{\mathbb{F}_{q}}$. Since $d=7$, this forces at least one of the irreducible components to also have multiplicity 1 and be defined over $\mathbb{F}_{q}$. This was a key insight used by Wooley [Woo08] to show that $p$-adic fields are $C_{2}(7)$ when q>883q>{883}. (resp. $C_{2}(11)$ when q>8053q>{8053}).
We will need a variant of Theorem 2.1 describing how often $X_{f}$ has linear factors when sliced by planes through a specified line $L$, under some hypotheses on $X_{f}\cap L$. Before stating and proving this variant in Corollary 2.7, we briefly describe a relaxation of Kaltofen’s factorization algorithm for $f(X,Y)$ over $\overline{k}$, which gives partial information when $f(X,0)$ is not squarefree.
2.1. A factorization algorithm
In this subsection, suppose $f(X,Y)\in k[X,Y]$ is monic in $X$ and has total degree $d$ (not necessarily homogeneous). Fix a positive integer $D<d$. The factorization algorithms found in [Kal95] and [CM06] underlying Theorem 2.1 request that $f(X,0)$ is squarefree. With slight modifications however, they still yield useful information when $f(X,0)$ is not squarefull, i.e. when considered over $\overline{k}$, $f(X,0)$ has at least one simple root. We recall the setup for the algorithm; see also [Kal95, §1], [CM06, §3.1].
Let $\alpha_{0}\in\overline{k}$ be a root of $f(X,0)$ such that $\frac{\partial f}{\partial x}(\alpha_{0},0)\neq 0$. Viewing $f$ as a univariate polynomial in $X$ over $k(\alpha_{0})[[Y]]$, since the partial derivative at $\alpha_{0}$ is a unit, we can lift $\alpha_{0}$ to $\widetilde{\alpha}\in k(\alpha_{0})[[Y]]$ satisfying $f(\widetilde{\alpha},Y)=0$. Set $\ell_{\max}=dD$, our desired order of approximation. Truncating, we produce $\alpha\in k(\alpha_{0})[Y]$ of $Y$-degree at most $\ell_{\max}$ satisfying
$f(\alpha,Y)$ $\equiv 0\pmod{Y^{\ell_{\max}+1}}$.
This may be accomplished explicitly via Newton’s method, as in [Kal95, CM06]. For $\mu=1,\ldots,D$ we write $\alpha^{\mu}$ as
$\alpha^{\mu}\equiv\sum_{r=0}^{\ell_{\max}}a_{\mu,r}Y^{r}\pmod{Y^{\ell_{\max}+1}}$,
where the $a_{\mu,r}\in k(\alpha_{0})$.
We now look for a polynomial $g\in\overline{k}[X,Y]$ of degree $m\leq D$ for which $g\mid f$ and $g(\alpha_{0},0)=0$. Here we will use $\ell=dm$ as our order of approximation. For each such $m$ we examine whether
(2.1) $\alpha^{m}+\sum_{\mu=0}^{m-1}h_{\mu}(Y)\alpha^{\mu}\equiv 0\pmod{Y^{\ell+1}}$
has a solution for $h_{\mu}(Y)\in\overline{k}[Y]$ with $\deg h_{\mu}\leq m-\mu$. Writing
$h_{\mu}(Y)=\sum_{\eta=0}^{m-\eta}u_{\mu,\eta}Y^{\eta}$
for $u_{\mu,\eta}\in\overline{k}$, we have that (2.1) holds if and only if the linear system
(2.2) $a_{m,r}+\sum_{\mu=0}^{m-1}\sum_{\eta=0}^{m-\mu}a_{\mu,r-\eta}u_{\mu,\eta}=0\text{ for }0\leq r\leq\ell$
has a solution in $\overline{k}$. Note that here we take the convention that $a_{\mu,r-\eta}=0$ when $\eta>r$.
Example 2.3 ($D=1$).
As an explicit example, we describe (2.1) and (2.2) in the special case of $D=1$. Here we take $m=1$, so we are looking to solve
$\alpha+h_{0}(Y)\equiv 0\pmod{Y^{d+1}}$,
for $h_{0}(Y)=u_{0,1}Y+u_{0,0}\in k(\alpha_{0})[Y]$. We have
$\alpha+h_{0}(Y)=a_{1,0}+u_{0,0}+(a_{1,1}+u_{0,1})Y+\sum_{r=2}^{d}a_{1,r}Y^{r},$
so a solution to (2.1) must have
$a_{1,0}+u_{0,0}$ $=0$
$a_{1,1}Y+u_{0,1}Y$ $=0$
$a_{1,2}Y^{2}$ $=0$
$\vdots$
$a_{1,d}Y^{d}$ $=0.$
Rearranging (2.2) into a matrix equation over $\overline{k}$, this looks like
$\begin{pmatrix}1&0\\ 0&1\\ 0&0\\ \vdots&\vdots\\ 0&0\end{pmatrix}\begin{pmatrix}u_{0,0}\\ u_{0,1}\end{pmatrix}=\begin{pmatrix}-a_{1,0}\\ -a_{1,1}\\ -a_{1,2}\\ \vdots\\ -a_{1,d}\end{pmatrix}$.
This system is insoluble for $d>1$, unless $a_{1,r}=0$ for $2\leq r\leq d$, i.e. $\alpha\equiv a_{1,0}+a_{1,1}Y\pmod{Y^{d+1}}$ all along.
When $D=1$, a solution to (2.2) is equivalent to $f$ possessing a linear factor vanishing on $(\alpha_{0},0)$. This is essentially a modified version of [CM06, Lemma 3.1].
Proposition 2.4.
Suppose $f(X,Y)\in k[X,Y]$ is a (not necessarily homogeneous) polynomial of total degree $d$ such that $f(X,0)$ is degree $d$ in $k[X]$ with a simple root $\alpha_{0}\in\overline{k}$.
Then $f$ factors over $\overline{k}$ as $f=g(X,Y)h(X,Y)$ for $\deg g=1$ satisfying $g(\alpha_{0},0)=0$ if and only if the linear system (2.2) has a solution in $k(\alpha_{0})$ for $m=1$.
Proof.
If (2.2) has a solution in $k(\alpha_{0})$ for $m=1$, then we have $\ell=d$ and there exists $g(X,Y)\in k(\alpha_{0})[X,Y]$, linear in $X$ and $Y$, such that $g(\alpha,Y)\equiv 0\pmod{Y^{\ell+1}}$. Consider $R(Y)=\mathrm{Res}_{X}(f,g)\in k(\alpha_{0})[Y]$, which is a polynomial in $Y$ of degree at most $d$. Since $f(\alpha,Y)\equiv 0\pmod{Y^{\ell+1}}$ as well, we have $R(Y)\equiv 0\pmod{Y^{\ell+1}}$, This forces $R=0$, so $f$ and $g$ have a nontrivial common factor in $k(\alpha_{0})[X,Y]$. But $g(X,Y)$ is already linear, so we have $g\mid f$ and $g(\alpha_{0},0)=0$.
Suppose $f(X,Y)=g(X,Y)h(X,Y)$ over $\overline{k}$ for $\deg g=1$ satisfying $g(\alpha_{0},0)=0$. Since $\alpha_{0}$ is a simple root of $f(X,0)$, we must in fact have $g$ defined over $k(\alpha_{0})$; otherwise, the conjugates of $g$ in $\overline{k}/k(\alpha_{0})$ would also vanish at $(\alpha_{0},0)$, producing a multiple root of $f(X,0)$ at $\alpha_{0}$. With $g,h$ defined over $k(\alpha_{0})$, we also have $h(\alpha_{0},0)\neq 0$, so $h(\alpha,Y)\not\equiv 0\pmod{Y}$. Since $f(\alpha,Y)\equiv 0\pmod{Y^{\ell_{\max}+1}}$, we have $g(\alpha,Y)\equiv 0\pmod{Y^{\ell_{\max}+1}}$. The coefficients of $g$ thus produce a solution to (2.2) in $k(\alpha_{0})$ for $m=1$. ∎
If $1<D\leq d-1$, we can still apply the strategy above to study the factorization of $f$. Given a solution to (2.2), we similarly produce $g$ of degree at most $D$ which has a common factor with $f$ and satisfies $g(\alpha_{0},0)=0$. However, when $D>1$, this does not guarantee that the common factor also vanishes on $(\alpha_{0},0)$.
For example, if $D=2$ and $f$ has a linear factor not vanishing on $(\alpha_{0},0)$, we could construct a solution to (2.2) with $m$ at most 2; we can no longer get away with degree 1, since we have asked for our factor $g$ to vanish on $(\alpha_{0},0)$. More generally for $D>1$, if (2.2) has no solutions in $\overline{k}$ for $m\leq D$, we may conclude that $f$ has no factors of degree at most $D$ vanishing on $(\alpha_{0},0)$, nor any factors of degree at most D1D-1. In particular, when D=d1D=d-1, this gives an irreducibility test which may be of independent interest; the proof is analogous to that of Proposition 2.4.
Proposition 2.5.
Suppose $f(X,Y)\in k[X,Y]$ is a (not necessarily homogeneous) polynomial of total degree $d$ such that $f(X,0)$ is degree $d$ in $k[X]$ with a simple root $\alpha_{0}\in\overline{k}$.
Then $f$ factors over $\overline{k}$ if and only if the linear system (2.2) has a solution in $k(\alpha_{0})$ for some $m<d$.
2.2. A genericity condition
Theorem 2.6.
Suppose $f(x_{1},\ldots,x_{n})$ is a (not necessarily homogeneous) polynomial of total degree $d$ over $k$, possessing no linear factors over $\overline{k}[x_{1},\ldots,x_{n}]$. Moreover, suppose there exist $(\mathbf{v}_{0},\mathbf{w}_{0})\in k^{2n-1}$ such that
$f_{\mathbf{v}_{0},\mathbf{w}_{0},\mathbf{z}}(X,0)=f(X+v_{1},w_{2}X+v_{2},\ldots,w_{n}X+v_{n})$
is a polynomial of degree $d$ in $k[X]$ with a simple root $\alpha_{0}\in\overline{k}$.
There exists $\Psi\in k(\alpha_{0})[Z_{2},\ldots,Z_{n}]$, depending on $\mathbf{v}_{0},\mathbf{w}_{0}$, of degree
$\deg\Psi\leq 3d-3$
such that $f_{\mathbf{v}_{0},\mathbf{w}_{0},\mathbf{z}}(X,Y)$ has no linear factor over $\overline{k}$ vanishing on $(\alpha_{0},0)$ whenever $\Psi(z_{2},\ldots,z_{n})\neq 0$.
Proof.
We follow the same lines as [Kal95, Theorem 5] and [CM06, Theorem 3.3]. Let
$\chi=\chi(X,Y,Z_{2},\ldots,Z_{n})=f(X+v_{1},w_{2}X+Z_{2}Y+v_{2},\ldots,w_{n}X+Z_{n}Y+v_{n}),$
considered in $\overline{k}[X,Y,Z_{2},\ldots,Z_{n}]$. We first establish that $\chi$ has no factor linear in $X,Z_{2},\ldots,Z_{n}$, reminiscent of [Kal95, Lemma 5].
Suppose $\chi=\chi_{1}\chi_{2}$ with $\chi_{1}$ linear in $X,Z_{2},\ldots,Z_{n}$ over $\overline{k}[Y]$. Viewing $\chi\in\overline{k}[Y,Z_{2},\ldots,Z_{n}][X]$, the leading term in $X$ is a constant in $k$ (depending on $\mathbf{w}_{0}$) times $X^{d}$, by hypothesis. Thus both $\chi_{1}$ and $\chi_{2}$ must have positive degree in $X$.
Under the ring map
$\overline{k}[X,Y,Z_{2},\ldots,Z_{n}]$ $\to\overline{k}(Y)[x_{1},\ldots,x_{n}]$
$X$ $\mapsto x_{1}-v_{1}$
$Z_{i}$ $\mapsto\frac{1}{Y}\left(x_{i}-w_{i}(x_{1}-v_{1})-v_{i}\right)\text{ for }2\leq i\leq n,$
we see that $\chi$ maps to the image of $f$ under the inclusion $\overline{k}[x_{1},\ldots,x_{n}]\to\overline{k}(Y)[x_{1},\ldots,x_{n}]$. The map also preserves the total degree in $X,Z_{2},\ldots,Z_{n}$ (it is invertible upon inverting $Y$ in the domain), so the image of $\chi_{1}$ remains linear in $x_{1},\ldots,x_{n}$ over $\overline{k}(Y)$.
Thus we have that the image of $f$ in $\overline{k}(Y)[x_{1},\ldots,x_{n}]$ has a linear factor in the image of $\chi_{1}$. Since $f$ does not depend on $Y$ in the first place, this implies $f$ has a linear factor in $\overline{k}[x_{1},\ldots,x_{n}]$, contradicting the hypothesis.
Now consider $\chi\in\overline{k}(Z_{2},\ldots,Z_{n})[X,Y]$. Since we have established that $\chi$ has no linear factors, let alone any containing $(\alpha_{0},0)$, by Proposition 2.4 the system (2.2) has no solutions in $\overline{k}(Z_{2},\ldots,Z_{n})(\alpha_{0})$ for $m=1$. Moreover, upon examination of the Newton iteration step [Kal95, CM06, Step N] we observe the system (2.2) has coefficients in $k(\alpha_{0})[Z_{2},\ldots,Z_{n}]$; the only denominators introduced are those involving $\frac{\partial f}{\partial x}(\alpha_{0},0)\in k(\alpha_{0})$.
Let $\Psi$ be a nonzero determinant of a $3\times 3$ minor of the augmented matrix of (2.2). If $\mathbf{z}\in k^{n-1}$ satisfies $\Psi(z_{2},\ldots,z_{n})\neq 0$ then the system (2.2) obtained by substituting $Z_{i}=z_{i}$ for $2\leq i\leq n$ is inconsistent. Thus by Proposition 2.4, $f_{\mathbf{v}_{0},\mathbf{w}_{0},\mathbf{z}}(X,Y)$ has no linear factors defined over $\overline{k}$ vanishing on $(\alpha_{0},0)$.
Following the degree computations for $\Psi_{1}^{(i)}$ in the proof of Theorem 2.1 (see also the proofs of [CM06, Theorem 3.3] and [Kal95, Theorem 5]), we find
$\deg\Psi\leq\sum_{j=0}^{2}\ell_{\max}-j=3d-3,$
completing the proof. ∎
Setting $k=\mathbb{F}_{q}$, and recognizing the number of $\mathbf{z}\in\mathbb{F}_{q}^{n-1}$ for which $\Psi(z_{2},\ldots,z_{n})=0$ is at most $(\deg\Psi)q^{n-2}$, we arrive at the following corollary.
Corollary 2.7.
Suppose $k=\mathbb{F}_{q}$ and $f$ satisfies the hypotheses of Theorem 2.6. Then there exist at most $(3d-3)q^{n-2}$ tuples $\mathbf{z}\in\mathbb{F}_{q}^{n-1}$ such that $f_{\mathbf{v}_{0},\mathbf{w}_{0},\mathbf{z}}$ has no linear factor over $\overline{\mathbb{F}_{q}}$ containing $(\alpha_{0},0)$.
2.3. Proof of Theorem 2.1
We conclude this section by giving a proof of Theorem 2.1. We follow the approach of Cafure and Matera [CM06, §3] and highlight the minor modification to the necessary order of approximation which leads to the modest improvement over [CM06, Corollaries 3.2, 3.4]; see also [Kal95].
Proof of Theorem 2.1.
Suppose that $f(X,Y)\in k[X,Y]$ is a monic polynomial of total degree $d$ over an arbitrary field $k$, such that $f(X,0)$ is squarefree. Cafure and Matera [CM06, §3.1] give an algorithm for testing whether such $f$ has a factor of degree at most $D<d$, which can be improved slightly in this case by setting $\ell_{\max}=dD$ (instead of $2dD$). The correctness of this modified algorithm is verified by the same argument as in the proof of [CM06, Lemma 3.1], similarly to Proposition 2.4 above. Following their notation, we have since the total degrees of $f$ and $g$ are $d$ and $m\leq D$, respectively, the resultant $\mathrm{Res}_{X}(f,g)\in\overline{k}[Y]$ has degree at most $dm$ in $Y$.
Returning to our absolutely irreducible polynomial $f\in\mathbb{F}_{q}[x_{1},\ldots,x_{n}]$, following [CM06, §3.2] and [Kal95, Lemmas 4, 5], there exists a polynomial $\Upsilon(v_{1},\ldots,v_{n},w_{2},\ldots,w_{n})$ of total degree at most $2d^{2}$ such that
$\chi(X,Y,Z_{2},\ldots,Z_{n})=f(X+v_{1},w_{2}X+Z_{2}Y+v_{2},\ldots,w_{n}X+Z_{n}Y+v_{n})$
is irreducible in $\overline{\mathbb{F}_{q}}[X,Y,Z_{2},\ldots,Z_{n}]$ with $\chi(X,0,Z_{2},\ldots,Z_{n})\in\overline{\mathbb{F}_{q}}[X]$ squarefree, whenever
$\Upsilon(v_{1},\ldots,v_{n},w_{2},\ldots,w_{n})\neq 0.$
The upshot is that (after possibly dividing by the leading coefficient in $X$) we can think of $\chi\in\overline{\mathbb{F}_{q}}(Z_{2},\ldots,Z_{n})[X,Y]$ and apply the aforementioned factorization algorithm. As in [CM06, Theorem 3.3], this produces a polynomial $\Psi_{D}^{(i)}\in\overline{\mathbb{F}_{q}}[Z_{2},\ldots,Z_{n}]$ such that for each root $\alpha_{i,0}\in\overline{\mathbb{F}_{q}}$ of $\chi(X,0,Z_{2},\ldots,Z_{n})$, if $\Psi_{D}^{(i)}(z_{2},\ldots,z_{n})\neq 0$, then $f_{\mathbf{v},\mathbf{w},\mathbf{z}}(X,Y)=\chi(X,Y,z_{2},\ldots,z_{n})$ has no factors of degree at most $D$ vanishing on $(\alpha_{i,0},0)$. Taking $\Psi_{D}=\prod\Psi_{D}^{(i)}$, we have that $f_{\mathbf{v},\mathbf{w},\mathbf{z}}$ has no factors of degree at most $D$ over $\overline{\mathbb{F}_{q}}$ whenever
$\Xi_{D}(v_{1},\ldots,v_{n},w_{2},\ldots,w_{n},z_{2},\ldots,z_{n})=\Upsilon(v_{1},\ldots,v_{n},w_{2},\ldots,w_{n})\Psi_{D}(z_{2},\ldots,z_{n})\neq 0.$
To estimate the degree of $\Psi_{D}$, and hence $\Xi_{D}$, the arguments of [CM06, Theorem 3.3] (see also [Kal95, Theorem 5]) remain unchanged, save for the updated value of $\ell_{\max}=dD$. We have
$\deg_{Z_{2},\ldots,Z_{n}}\Psi_{D}^{(i)}$ $\leq\sum_{j=0}^{(D+1)(D+2)/2-1}\ell_{\max}-j$
$=\frac{1}{8}\left(-D^{4}+4D^{3}d-6D^{3}+12D^{2}d-11D^{2}+8Dd-6D\right)$
Multiplying this bound by $d$ yields an upper bound for $\deg_{Z_{2},\ldots,Z_{n}}\Psi_{D}$, and adding $2d^{2}$ to account for $\deg\Upsilon$ gives
$\deg\Xi_{D}\leq\frac{d}{8}\left(-D^{4}+4D^{3}d-6D^{3}+12D^{2}d-11D^{2}+8Dd-6D+16d\right).$
Combining this with the well known fact (see e.g. [CM06, Lemma 2.1])
$\{(\mathbf{v},\mathbf{w},\mathbf{z})\in\mathbb{F}_{q}^{3n-2}:\Xi_{D}(\mathbf{v},\mathbf{w},\mathbf{z})=0\}\leq(\deg\Xi_{D})q^{3n-3}$
gives the stated result (ii).
The result (i) follows from recognizing that we need only that $\Xi=\Upsilon\Psi_{d-1}^{(1)}\neq 0$ to ensure that $f_{\mathbf{v},\mathbf{w},\mathbf{z}}$ is absolutely irreducible, since at least one factor of $f_{\mathbf{v},\mathbf{w},\mathbf{z}}$ must vanish on $(\alpha_{1,0},0)$. The corresponding degree bound yields the stated result. ∎
3. Facts about degree 7 forms over $\mathbb{F}_{q}$
Throughout this section, we take $f(x_{0},\ldots,x_{m})$ to be a nonzero homogeneous form of degree $d$ over a finite field $\mathbb{F}_{q}$ and denote by $X_{f}\subset\mathbb{P}^{m}$ the degree $d$ hypersurface cut out by $f=0$. We briefly recall some generalities for degree $d$ forms, before specializing to degree $d=7$.
Lemma 3.1.
Suppose $q\geq{d}$ and $f$ is a nonzero degree $d$ form in $m+1\geq 2$ variables. Then $X_{f}(\mathbb{F}_{q})\subsetneq\mathbb{P}^{m}(\mathbb{F}_{q})$.
Proof.
Suppose $m=1$, i.e. $f$ is a binary form of degree $d$ that vanishes on all $q+1$ points of $\mathbb{P}^{1}(\mathbb{F}_{q})$. Then $x_{0}^{q}x_{1}-x_{0}x_{1}^{q}\mid f$, so we have $\deg f\geq q+1>d$, a contradiction.
We proceed by induction on $m$. If $f$ vanishes on all points in $\mathbb{P}^{m}(\mathbb{F}_{q})$, then $f(x_{0},\ldots,x_{m-1},0)\in\mathbb{F}_{q}[x_{0},\ldots,x_{m-1}]$ is a (possibly now zero) degree $d$ form in $m$ variables that vanishes on all $\mathbb{F}_{q}$-points of $\mathbb{P}^{m-1}$. By the induction hypothesis, it must indeed be the zero form, i.e. $x_{m}\mid f$.
Repeating for all possible distinct (up to scaling) linear factors of $f$, of which there are more than $d$, we have found too many linear factors of $f$, forcing $f=0$, a contradiction. ∎
A consequence of Lemma 3.1 is that when $q\geq{d}$ and $H\subset\mathbb{P}^{m}$ is a hyperplane defined over $\mathbb{F}_{q}$, $X_{f}$ vanishes on all $\mathbb{F}_{q}$-points of $H$ if and only if $H\subseteq X_{f}$, i.e. $f$ has a linear factor which corresponds to $H$. Moving forward, we will often conflate these notions.
Let us set a bit more notation. We refer to subvarieties $U\subset\mathbb{P}^{m}$ cut out by the vanishing of a collection of linear forms as linear subvarieties. Given such $U$ defined over $\mathbb{F}_{q}$ of dimension $s$, we use $f|_{U}$ to refer to the form cutting out $X_{f}\cap U$. If $U\subset X_{f}$, then $f|_{U}=0$; otherwise $f|_{U}$ is a nonzero degree $d$ form in $s+1$ variables. Note that $f|_{U}$ is only well defined up to choice of coordinates on $U$ and scaling, but this will not be of concern.
Given linear subvarieties $U,V\subseteq\mathbb{P}^{m}$, their span, denoted $\langle U,V\rangle$, is the minimal linear subvariety of $\mathbb{P}^{m}$ containing $U$ and $V$. In particular, when $U$ and $V$ are $\mathbb{F}_{q}$-points, $\langle U,V\rangle$ denotes the line between them.
For the remainder of the section, $f=f(x_{0},\ldots,x_{m})$ denotes a nonzero degree 7 form in $m+1$ variables over $\mathbb{F}_{q}$. We assume $m>1$.
Lemma 3.2.
Let $q\geq{7}$. Suppose $U,V\subset X_{f}$ are distinct linear subvarieties of dimension $s\geq 0$, defined over $\mathbb{F}_{q}$, such that $\dim(U\cap V)=s-1$. [^1] Throughout, we take the convention that an empty intersection has dimension -1 and codimension $s+1$ in a linear variety of dimension $s$. Then one of the following holds.
  1. (i)
    $f|_{\langle U,V\rangle}$ has an absolutely irreducible factor of multiplicity 1 over $\mathbb{F}_{q}$ of degree at most 3.
  2. (ii)
    $(X_{f}\cap\langle U,V\rangle)(\mathbb{F}_{q})=U\cup V\cup W$ for a linear subvariety $W\subset X_{f}$ (possibly empty) of dimension at most $s$ and defined over $\mathbb{F}_{q}$.
  3. (iii)
    $f|_{\langle U,V\rangle}=0$, i.e. $\langle U,V\rangle\subset X_{f}$.
Proof.
Assume $f|_{\langle U,V\rangle}\neq 0$, i.e. (iii) does not hold. Then $f|_{\langle U,V\rangle}$ is a (nonzero) degree $d$ form in $s+2$ variables, coming from $\langle U,V\rangle\simeq\mathbb{P}^{s+1}$.
$f|_{\langle U,V\rangle}$ has two distinct linear factors $\ell_{U},\ell_{V}$ coming from the fact that it vanishes on $U,V$, since both are codimension 1 in $\langle U,V\rangle$. If either factor has multiplicity 1, then we are in case (i). Otherwise, we write
$f|_{\langle U,V\rangle}=\ell_{U}^{m_{U}}\ell_{V}^{m_{V}}g$
for $m_{U},m_{V}\geq 2$ and $g$ a form of degree $7-m_{U}-m_{V}\leq 3$ over $\mathbb{F}_{q}$.
If $g$ is irreducible over $\overline{\mathbb{F}_{q}}$, we are again in case (i), so assume $g$ is reducible over $\overline{\mathbb{F}_{q}}$. The only way for it to fail to have an irreducible factor of multiplicity 1 over $\mathbb{F}_{q}$ is for $g=\ell_{W}^{m_{W}}$ for $m_{W}=7-m_{U}-m_{V}$, or for $g$ to be irreducible over $\mathbb{F}_{q}$ but factor into conjugate linear forms over $\overline{\mathbb{F}_{q}}$.
In the former case, we have that $f|_{\langle U,V\rangle}$ vanishes on $W\colon\ell_{W}=0$ of dimension $s$. In the latter, we set $W$ to be the (possibly empty) intersection of the linear subspaces over $\overline{\mathbb{F}_{q}}$ cut out by the linear factors of $g$ defined over $\overline{\mathbb{F}_{q}}$. Both of these fall under case (ii). ∎
We would like a version of Lemma 3.2 where (i) implies $X_{f}(\mathbb{F}_{q})^{\mathrm{sm}}\neq\emptyset$. To achieve this, we need a lower bound for the number of smooth $\mathbb{F}_{q}$-points on a degree $d$ plane curve.
Lemma 3.3.
Let $C\subset\mathbb{P}^{2}$ be a plane curve defined over $\mathbb{F}_{q}$ of degree $d$. Suppose $C^{\prime}\subset C$ is a geometrically integral component, defined over $\mathbb{F}_{q}$, of multiplicity 1 and degree $d^{\prime}$. Denote by $g^{\prime}$ the geometric genus of the normalization of $C^{\prime}$. Then we have
(3.1) $C(\mathbb{F}_{q})^{\mathrm{sm}}\geq q+1-g^{\prime}\lfloor 2\sqrt{q}\rfloor-(d^{\prime}-1)(d^{\prime}-2)+2g^{\prime}-(d-d^{\prime})d^{\prime}.$
Proof.
We begin by counting points on $C^{\prime}$. The usual Hasse–Weil bounds can be modified to allow for $C^{\prime}$ to be singular, yielding
(3.2) $C^{\prime}(\mathbb{F}_{q})\geq q+1-g^{\prime}\lfloor 2\sqrt{q}\rfloor-\frac{1}{2}(d^{\prime}-1)(d^{\prime}-2)+g^{\prime}$
(see e.g. [LY94, Corollary 1]). There are at most $\frac{1}{2}(d^{\prime}-1)(d^{\prime}-2)-g^{\prime}$ singular $\mathbb{F}_{q}$-points on $C^{\prime}$, so from (3.2) we deduce
$\#C^{\prime}(\mathbb{F}_{q})^{\mathrm{sm}}$ $=\#C^{\prime}(\mathbb{F}_{q})-\#C^{\prime}(\mathbb{F}_{q})^{\mathrm{sing}}$
(3.3) $\geq q+1-g^{\prime}\lfloor 2\sqrt{q}\rfloor-(d^{\prime}-1)(d^{\prime}-2)+2g^{\prime}.$
A point in $C^{\prime}(\mathbb{F}_{q})^{\mathrm{sm}}$ is a smooth point of $C$ if it does not also lie on $C-C^{\prime}$. By Bezout’s theorem, there are at most $d^{\prime}(d-d^{\prime})$ such intersection points. Thus (3.1) follows from subtracting this from (3). ∎
Remark 3.4.
For constant $d^{\prime}$ and q>2q>{2}, the lower bound (3.1) is monotonically decreasing in $g^{\prime}$. In other words, for $C^{\prime}$ of fixed degree, the $q$-threshhold where the lower bound becomes positive is largest when $g^{\prime}$ is as large as possible.
Lemma 3.5.
Let q>39q>{39}. Suppose $U,V\subset X_{f}$ are as in Lemma 3.2 with $s\geq 1$. Then one of the following holds.
  1. (i)
    $X_{f}(\mathbb{F}_{q})^{\mathrm{sm}}\neq\emptyset$.
  2. (ii)
    $(X_{f}\cap\langle U,V\rangle)(\mathbb{F}_{q})=U\cup V\cup W$ for $W\subset X_{f}$ linear (possibly empty) of dimension at most $1$ and defined over $\mathbb{F}_{q}$.
  3. (iii)
    $f|_{\langle U,V\rangle}=0$, i.e. $\langle U,V\rangle\subset X_{f}$.
Proof.
By Lemma 3.2, we have either (ii) or (iii) hold, or that $f|_{\langle U,V\rangle}$ has a simple irreducible factor $g$ of degree at most $3$. In this case, we argue that $X_{f}(\mathbb{F}_{q})^{\mathrm{sm}}\neq\emptyset$.
Suppose $\deg g=1$. Then $X_{f}$ has a linear irreducible component $L\simeq\mathbb{P}^{s}$ of multiplicity 1 defined over $\mathbb{F}_{q}$. In particular, $L(\mathbb{F}_{q})=L(\mathbb{F}_{q})^{\mathrm{sm}}=\frac{q^{s+1}-1}{q-1}$. Since $(X_{f}\cap\langle U,V\rangle)-L\subset\mathbb{P}^{s+1}$ is a degree 6 hypersurface, its intersection with $L$ is also a degree 6 hypersurface in $L\simeq\mathbb{P}^{s}$. Such a hypersurface has at most $6\frac{q^{s}-1}{q-1}$ $\mathbb{F}_{q}$-points, so we can find an $\mathbb{F}_{q}$-point in $L(\mathbb{F}_{q})-((X_{f}\cap\langle U,V\rangle)-L)(\mathbb{F}_{q})$, which is smooth in $X_{f}$.
Suppose $f|_{\langle U,V\rangle}$ has no linear components of multiplicity 1. Then the irreducible components of $X_{f}\cap\langle U,V\rangle$ are $U$, $V$, and one cut out by $g=0$, the latter of which is geometrically integral of degree 2 or 3.
Consider first the case of $s=1$, where $f|_{\langle U,V\rangle}$ cuts out a degree 7 plane curve $C$ with a geometrically irreducible component $C^{\prime}$ defined over $\mathbb{F}_{q}$ of degree 2 or 3. Applying Lemma 3.3 with $d=7$, $d^{\prime}=2,3$ and $0\leq g^{\prime}\leq\frac{(d^{\prime}-1)(d^{\prime}-2)}{2}$, we conclude q>19q>{19} suffices to ensure $C_{f}(\mathbb{F}_{q})^{\mathrm{sm}}>0$, hence $X_{f}(\mathbb{F}_{q})^{\mathrm{sm}}\neq\emptyset$.
Suppose now that $s>1$. We reduce to the $s=1$ case by an application of Theorem 2.1. After change of coordinates we may assume $U\colon x_{0}=0$ and $V\colon x_{1}=0$ and write
$f=x_{0}^{m_{U}}x_{1}^{m_{V}}g(x_{0},\ldots,x_{m})$
for $m_{U},m_{V}\geq 2$ and $g$ absolutely irreducible of degree 2 or 3. Dehomogenizing by setting $x_{0}=1$ and applying Theorem 2.1(i), when q>39q>{39} there exist $(\mathbf{v},\mathbf{w},\mathbf{z})$ such that $g_{\mathbf{v},\mathbf{w},\mathbf{z}}$ is absolutely irreducible. Now we have
$f_{\mathbf{v},\mathbf{w},\mathbf{z}}(X,Y)=(X-v_{1})^{m_{V}}g_{\mathbf{v},\mathbf{w},\mathbf{z}}(X,Y).$
Applying the same argument as in the $s=1$ case above (rehomogenizing if necessary), we see that q>39q>{39} is more than sufficient to guarantee the existence of a smooth $\mathbb{F}_{q}$-solution to $f_{\mathbf{v},\mathbf{w},\mathbf{z}}=0$, and hence $X_{f}(\mathbb{F}_{q})^{\mathrm{sm}}\neq\emptyset$. ∎
We next give an intermediate result in a similar spirit to [LY96, Lemma 5.3], which will be useful later.
Lemma 3.6.
Let q>7q>{7}. Suppose $U\subset X_{f}$ is a linear subvariety of positive dimension defined over $\mathbb{F}_{q}$, and $v\in X_{f}(\mathbb{F}_{q})-U$. If there exist distinct codimension 1 linear subvarieties $U_{1}^{(1)},U_{2}^{(1)},U_{3}^{(1)}\subset U$ such that
$\langle U_{i}^{(1)},v\rangle\subset X_{f}\text{ for }i=1,2,3,$
then either $X_{f}(\mathbb{F}_{q})^{\mathrm{sm}}\neq\emptyset$ or $\langle U,v\rangle\subset X_{f}$.
Proof.
If $f|_{\langle U,v\rangle}\neq 0$ then $f|_{\langle U,v\rangle}$ has at least four distinct linear factors cutting out the codimension 1 linear subvarieties $U,\langle U_{1}^{(1)},v\rangle,\langle U_{2}^{(1)},v\rangle,\langle U_{3}^{(1)},v\rangle\subset\langle U,v\rangle$. One of these components must have multiplicity 1, so by the argument in the proof of Lemma 3.5 we have $X_{f}(\mathbb{F}_{q})^{\mathrm{sm}}\neq\emptyset$. ∎
We conclude this section by examining the situation of a plane curve of degree 7 over $\mathbb{F}_{q}$ which has at least three non-colinear $\mathbb{F}_{q}$-points and no linear components defined over $\mathbb{F}_{q}$. Unfortunately, we cannot guarantee the existence of a smooth $\mathbb{F}_{q}$-point, even when $q$ is sufficiently large. We can, however, find either a smooth $\mathbb{F}_{q}$-point or a line spanned by (singular) $\mathbb{F}_{q}$-points which elsewhere meets the curve at a nonsingular geometric point.
Lemma 3.7.
Let q>591q>{591} and suppose $f\in\mathbb{F}_{q}[x_{0},x_{1},x_{2}]$ is a nonzero ternary degree 7 form with no linear factors defined over $\mathbb{F}_{q}$. Suppose further that there exist $u,v,w\in X_{f}(\mathbb{F}_{q})$ such that $\langle u,v,w\rangle=\mathbb{P}^{2}$. Then either $X_{f}(\mathbb{F}_{q})^{\mathrm{sm}}\neq\emptyset$ or there exist $u_{1},u_{2}\in X_{f}(\mathbb{F}_{q})$ such that $\langle u_{1},u_{2}\rangle\not\subset X_{f}$ and $X_{f}\cap\langle u_{1},u_{2}\rangle$ contains a geometric point of multiplicity 1.
Proof.
If any of $u,v,w\in X_{f}(\mathbb{F}_{q})^{\mathrm{sm}}$, then we are done, so we assume they are all singular. Since $f$ has no linear factors defined over $\mathbb{F}_{q}$, the possible factorization types of $f$ over $\mathbb{F}_{q}$ are $(7),(5,2),(4,3),(3,2,2),(3,2^{2})$.
Suppose one of these factors of multiplicity 1 remains irreducible over $\overline{\mathbb{F}_{q}}$. Let $C^{\prime}\subset X_{f}$ denote the corresponding irreducible component curve of degree $d^{\prime}$, defined over $\mathbb{F}_{q}$ and absolutely irreducible. If $d^{\prime}=7$ then the geometric genus of the normalization of $X_{f}$ is at most 12, owing to the presence of the singular points $u,v,w$. Lemma 3.3 gives that q>591q>{591} guarantees $X_{f}(\mathbb{F}_{q})^{\mathrm{sm}}\neq\emptyset$. If $d^{\prime}<7$ then the hypothesis q>591q>{591} is sufficient to show $X_{f}(\mathbb{F}_{q})^{\mathrm{sm}}\neq\emptyset$ by Lemma 3.3.
Assume now that all of the $\mathbb{F}_{q}$-factors of $f$ with multiplicity 1 are reducible over $\overline{\mathbb{F}_{q}}$. In the case of factorization type (7), the only possibility is that $f$ factors into 7 linear forms conjugate over $\mathbb{F}_{q^{7}}$. Such $f$ can vanish on at most one $\mathbb{F}_{q}$-point, contradicting our hypothesis that $u,v,w\in X_{f}(\mathbb{F}_{q})$. Similarly, in the case of $(5,2)$ we have $f$ can vanish on at most two $\mathbb{F}_{q}$-points, again contradicting the hypothesis.
In the remaining factorization types we see that $f$ has a cubic factor which is geometrically reducible, necessarily factoring as a product of three linear forms conjugate over $\mathbb{F}_{q^{3}}$. Let $L_{1},L_{2},L_{3}$ be the corresponding lines in the plane defined over $\mathbb{F}_{q^{3}}$.
Suppose $L_{1},L_{2},L_{3}$ meet at a common point $x\in X_{f}(\mathbb{F}_{q})$. Since the points $u,v,w$ are not collinear, we must have that $x$ does not lie on at least one of the lines $\langle u,v\rangle,\langle u,w\rangle,\langle v,w\rangle$. Without loss of generality, assume $x\notin\langle u,v\rangle$. Thus $f|_{\langle u,v\rangle}$ vanishes to order 2 on $u$ and $v$ and order 1 on the three distinct $\mathbb{F}_{q^{3}}$-points $\langle u,v\rangle\cap L_{i}$, so we may take $u_{1}=u$ and $u_{2}=v$.
If $L_{1},L_{2},L_{3}$ do not meet at a common point, then they contain no $\mathbb{F}_{q}$-points. Again considering the line $\langle u,v\rangle$, we have that at most two of of the $\mathbb{F}_{q^{3}}$-points $\langle u,v\rangle\cap L_{i}$ can coincide. Thus $f|_{\langle u,v\rangle}$ vanishes to order 2 on $u$ and $v$ and order 1 on at least one of the $\langle u,v\rangle\cap L_{i}$, so we may take $u_{1}=u$ and $u_{2}=v$. ∎
4. Reduced forms and consequences
We recall the setting of Artin’s conjecture for degree $d$ forms. $K/\mathbb{Q}_{p}$ is a $p$-adic field with valuation ring $\mathcal{O}_{K}$, uniformizer $\pi$, normalized discrete valuation $v$, and residue field $\mathcal{O}_{K}/(\pi)\simeq\mathbb{F}_{q}$. We have a homogeneous form $F\in K[x_{0},\ldots,x_{n}]$ of degree $d$ for $n\geq d^{2}$. After rescaling, we may assume $F\in\mathcal{O}_{K}[x_{0},\ldots,x_{n}]$ such that $\overline{F}\neq 0$ in $\mathbb{F}_{q}[x_{0},\ldots,x_{n}]$. We will use the lowercase $f=\overline{F}$ to denote the reduction to the residue field $\mathbb{F}_{q}$, and $X_{f}/\mathbb{F}_{q}$ for the special fiber of $X_{F}$ over $\mathcal{O}_{K}$.
By Hensel’s lemma, if $X_{f}(\mathbb{F}_{q})^{\mathrm{sm}}\neq\emptyset$, then $X_{F}(K)\neq\emptyset$. Therefore our goal will be to show that $X_{f}$ has a smooth $\mathbb{F}_{q}$-point when $q$ is sufficiently large. A key strategy in achieving this was introduced by Birch and Lewis [BL59] in the degree $d=5$ case, then later refined by Laxton and Lewis [LL65] to apply in degrees $d=7,11$. They define a notion of reduced forms which we recall now before establishing the relevant consequences. Both [LL65] and [LY96, §4] are useful references.
4.1. Reduced forms
With $F$ as above, let
$I(F)=\mathrm{Res}\left(\frac{\partial F}{\partial x_{0}},\ldots,\frac{\partial F}{\partial x_{n}}\right)\in\mathcal{O}_{K}.$
By the general theory of (multivariate) resultants, $I(F)$ is homogeneous of degree $(d-1)^{n}$ in the coefficients of the partial derivatives of $F$. We also have [LL65, Lemma 5]
$I(aF(T\mathbf{x}))=a^{(n+1)(d-1)^{n}}(\det T)^{d(d-1)^{n}}I(F)$
for $a\in K^{\times}$ and $T\in\operatorname{GL}_{n+1}(K)$ acting on the variables $(x_{0},\ldots,x_{n})$ by linear change of coordinates.
Definition 4.1 (equivalent forms).
Two degree $d$ forms $F,F^{\prime}\in\mathcal{O}_{K}[x_{0},\ldots,x_{n}]$ are equivalent if there exists $a\in K^{\times}$ and $T\in\operatorname{GL}_{n+1}(K)$ such that $F^{\prime}=aF(T\mathbf{x})$.
Definition 4.2 (reduced form).
We say a degree $d$ form $F\in\mathcal{O}_{K}[x_{0},\ldots,x_{n}]$ is reduced if $I(F)\neq 0$ and $v(I(F))\leq v(I(aF(T\mathbf{x})))$ for all $a\in K^{\times}$ and $T\in\operatorname{GL}_{n+1}(K)$ such that $aF(T\mathbf{x})$ has coefficients in $\mathcal{O}_{K}$. That is, $I(F)$ has minimal valuation among equivalent forms with coefficients in $\mathcal{O}_{K}$.
Proposition 4.3 ([LL65, LY96]).
We have the following facts about reduced forms.
  1. (i)
    If for all reduced degree $d$ forms $F\in\mathcal{O}_{K}[x_{0},\ldots,x_{n}]$ we have $X_{F}(K)\neq\emptyset$ then $X_{F}(K)\neq\emptyset$ for all degree $d$ forms $F\in\mathcal{O}_{K}[x_{0},\ldots,x_{n}]$, i.e. $K$ is $C_{2}(d)$.
  2. (ii)
    If $n\geq d^{2}$ and $F$ is a reduced degree $d$ form, then its reduction $f=\overline{F}$ has no linear factors defined over $\overline{\mathbb{F}_{q}}$.
  3. (iii)
    Let $q\geq{d}$. If $n\geq d^{2}$, $F$ is a reduced degree $d$ form, and $f$ vanishes on an $s$-dimensional linear subvariety $U\simeq\mathbb{P}^{s}$ defined over $\mathbb{F}_{q}$, then
    $X_{f}(\mathbb{F}_{q})\geq\frac{q^{s+2}-1}{q-1}.$
Proof.
It is established in [LL65, Corollary] that to show every $F$ has a $K$-solution, it suffices to show this only for forms with $I(F)\neq 0$. After possibly replacing $F$ by an equivalent form, we may assume $I(F)$ has minimal valuation, i.e. $F$ is reduced. This does not change solubility, so it suffices to check $X_{F}(K)\neq 0$ for reduced forms, giving (i).
(ii) is [LL65, Lemma 9], while (iii) is [LY96, Corollary 4.4]. In both cases, the idea is to use the fact that $F$ is a reduced form to bound the order of $f$ (or its factors), i.e. to bound the minimal number of variables appearing nontrivially in $f$, up to change of coordinates. To prove (ii), one understands that linear factors have small order which is at odds with these lower bounds; for (iii) an order estimate is combined with Warning’s estimate for $X_{f}(\mathbb{F}_{q})$ (see e.g. [LY96, Lemma 3.1]). ∎
4.2. Consequences
Drawing inspiration from the work of Leep and Yeomans in the quintic case [LY96], we seek to exploit what we know about reduced forms to produce a nonsingular point of $X_{f}$ which can be lifted to $X_{F}(K)$.
Our immediate goal, accomplished below in Proposition 4.5, is show that either $X_{f}(\mathbb{F}_{q})^{\mathrm{sm}}\neq\emptyset$ or we can produce two points $u,v\in X_{f}(\mathbb{F}_{q})$ such that $f$ possesses a nonsingular geometric point on the line between them. Later, in the proof of Theorem C, we will use this in conjunction with Theorem 2.6 to produce a planar slice of $X_{f}$ with at least two (possibly singular) $\mathbb{F}_{q}$-points. The presence of these two $\mathbb{F}_{q}$-points is key to improving on [Woo08, Theorem 1].
As in [LY96], we will make use of a linear subvariety $U\subset X_{f}$ of maximal dimension. In the following lemma, we show that not only can we find points $v,w\in(X_{f}-U)(\mathbb{F}_{q})$, but we can ensure the line $\langle v,w\rangle$ is skew to $U$. This is accomplished by strengthening an argument presented in [LY96, Proof of Proposition 5.4].
Lemma 4.4.
Let q>7q>{7} and suppose $F\in\mathcal{O}_{K}[x_{0},\ldots,x_{n}]$ is a reduced degree 7 form for $n\geq 49$. Suppose $U\subset X_{f}$ is a linear subvariety defined over $\mathbb{F}_{q}$ of maximal dimension. Then either $X_{f}(F_{q})^{\mathrm{sm}}\neq\emptyset$ or there exist $v,w\in X_{f}(\mathbb{F}_{q})$ such that
$\langle v,w\rangle\not\subset X_{f}\quad\text{and}\quad\langle v,w\rangle\cap U=\emptyset.$
Proof.
We have $U(\mathbb{F}_{q})=\frac{q^{s+1}-1}{q-1}<X_{f}(\mathbb{F}_{q})$ by Proposition 4.3(iii), so there exists $v\in X_{f}(\mathbb{F}_{q})-U(\mathbb{F}_{q})$. Moreover, we have $\langle U,v\rangle\not\subset X_{f}$ by maximality of $U$, and since q>7q>{7} an application of Lemma 3.1 reveals that $f$ does not vanish on all $\mathbb{F}_{q}$-points of $\langle U,v\rangle$. Since $\langle U,v\rangle(\mathbb{F}_{q})=\frac{q^{s+2}-1}{q-1}$, Proposition 4.3(iii) again ensures there exists a point $w\in(X_{f}-\langle U,v\rangle)(\mathbb{F}_{q})$.
We now have $\langle v,w\rangle\cap U=\emptyset$. If $\langle v,w\rangle\not\subset X_{f}$ then we are done. Suppose instead that for all $v,w\in(X_{f}-U)(\mathbb{F}_{q})$ such that $\langle v,w\rangle\cap U=\emptyset$, we have $\langle v,w\rangle\subset X_{f}$. Let $W\subset X_{f}$ be a linear subvariety over $\mathbb{F}_{q}$ of maximal dimension such that $W\cap U\subset W$ has codimension at least 2. Such $W$ exists because earlier we found a line $\langle v,w\rangle\subset X_{f}$ skew to UU. [^2] By convention, for $W$ a linear variety of dimension $s$, the codimension of $\emptyset\subset W$ is $s+1$.
We claim that $W$ contains $(X_{f}-U)(\mathbb{F}_{q})$. Suppose to the contrary that $x\in(X_{f}-U-W)(\mathbb{F}_{q})$. Since $U\cap W\subset W$ has codimension at least 2, we have that $\langle U,x\rangle\not\supset\langle W,x\rangle$.
Let $y\in\langle W,x\rangle(\mathbb{F}_{q})$. If $y\notin\langle U,x\rangle\cap\langle W,x\rangle$, then $\langle x,y\rangle\cap U=\emptyset$. Since $y$ is contained on a line between $x$ and a point of $W$, by assumption we have $\langle x,y\rangle\subset X_{f}$, and in particular, $y\in X_{f}(\mathbb{F}_{q})$. Thus $f$ vanishes on all of $\langle W,x\rangle(\mathbb{F}_{q})$, save possibly for a a subspace of codimension at least 1 arising from $\langle U,x\rangle\cap\langle W,x\rangle$. But then on $\langle W,x\rangle$, we have that there exists a nonzero linear form $\ell$ such that $\ell f|_{\langle W,x\rangle}$ vanishes on all points; by Lemma 3.1, $\ell f|_{\langle W,x\rangle}=0$. Since $\ell\neq 0$, we have $f|_{\langle W,x\rangle}=0$, i.e. $\langle W,x\rangle\subset X_{f}$.
Note that the dimension of $U\cap\langle W,x\rangle$ is at most one more than that of $U\cap W$, so its codimension is still at most 2 in $\langle W,x\rangle$. However, this violates the maximality of $W$, yielding a contradiction. We conclude that $(X_{f}-U)(\mathbb{F}_{q})\subset W(\mathbb{F}_{q})$, as claimed.
Now we have $X_{f}(\mathbb{F}_{q})=U(\mathbb{F}_{q})\cup W(\mathbb{F}_{q})$. However, counting $\mathbb{F}_{q}$-points we see
$X_{f}(\mathbb{F}_{q})\leq\#U(\mathbb{F}_{q})+\#W(\mathbb{F}_{q})\leq 2\frac{q^{s+1}-1}{q-1}<X_{f}(\mathbb{F}_{q}),$
by Proposition 4.3(iii), yielding another contradiction. Thus there exist $v,w\in X_{f}(\mathbb{F}_{q})-U(\mathbb{F}_{q})$ satisfying the desired hypotheses. ∎
Proposition 4.5.
Let q>591q>{591} and suppose $F\in\mathcal{O}_{K}[x_{0},\ldots,x_{n}]$ is a reduced degree 7 form for $n\geq 49$. Then either $X_{f}(\mathbb{F}_{q})^{\mathrm{sm}}\neq\emptyset$ or there exist $u_{1},u_{2}\in X_{f}(\mathbb{F}_{q})$ such that $\langle u_{1},u_{2}\rangle\not\subset X_{f}$ and $X_{f}\cap\langle u_{1},u_{2}\rangle$ contains a geometric point of multiplicity 1.
Proof.
Let $U\subset X_{f}$ be a linear subvariety defined over $\mathbb{F}_{q}$ of maximal dimension, denoted by $s$. We have $s\geq 0$ by Proposition 4.3(iii). We treat the $s=0$, $s=1$, and $s\geq 2$ cases separately.
Case $s=0$. If $s=0$, then by Proposition 4.3(iii) we have $X_{f}(\mathbb{F}_{q})\geq q+1>3$, so there exist at least four $\mathbb{F}_{q}$-points on $X_{f}$. If any four of them are colinear, then at least one has multiplicity 1, so we have $X_{f}(\mathbb{F}_{q})^{\mathrm{sm}}\neq\emptyset$. Thus there exist three points $u,v,w\in X_{f}(\mathbb{F}_{q})$ such that $\langle u,v,w\rangle\simeq\mathbb{P}^{2}$. Applying Lemma 3.7 to $f|_{\langle u,v,w\rangle}$, we have the desired result.
Case $s=1$. Let $U\subset X_{f}$ be a line defined over $\mathbb{F}_{q}$. By Lemma 4.4 there exist $v,w\in X_{f}(\mathbb{F}_{q})$ such that the line $\langle v,w\rangle$ is skew to $U$ and not contained in $X_{f}$. If for any point $u\in U(\mathbb{F}_{q})$ we have that $f$ vanishes on no lines in the plane $\langle u,v,w\rangle$, then we may apply Lemma 3.7, as in the $s=0$ case.
Assume to the contrary that for all $u\in U(\mathbb{F}_{q})$ we have that $f$ vanishes on some line in $\langle u,v,w\rangle$. By Lemma 3.6, either $X_{f}(\mathbb{F}_{q})^{\mathrm{sm}}\neq\emptyset$ or there are at most two points $u\in U(\mathbb{F}_{q})$ such that $\langle u,v\rangle\subset X_{f}$, since $\langle U,v\rangle\not\subset X_{f}$ in this case as $s=1$. Similarly, there are at most two $u\in U(\mathbb{F}_{q})$ for which $\langle u,w\rangle\subset X_{f}$, and if there is a third point $x\in(\langle v,w\rangle\cap X_{f})(\mathbb{F}_{q})$, at most two $u\in U(\mathbb{F}_{q})$ for which $\langle u,x\rangle\subset X_{f}$.
Thus for all but at most 6 points $u\in U(\mathbb{F}_{q})$, we have that $f$ does not vanish on any lines between $u$ and a point on $(X_{f}\cap\langle v,w\rangle)(\mathbb{F}_{q})\supset\{v,w\}$. Instead, $f$ must vanish on a line containing either $v$, $w$, or the potential third point $x$ on $(X_{f}\cap\langle v,w\rangle)(\mathbb{F}_{q})$. When q>17q>{17}, this guarantees the existence of at least five points $u_{1},\ldots,u_{5}\in U(\mathbb{F}_{q})$ such that $f$ vanishes on a line $V_{i}\subset\langle u_{i},v,w\rangle$ through a common point $v$, $w$, or $x$. Without loss of generality, assume $v\in V_{i}$ for $i=1,\ldots,5$. Note that $V_{i}\cap U=\emptyset$; if not we find that $w\in\langle U,v\rangle$, contradicting our choice of $\langle v,w\rangle$ skew to $U$.
Suppose for $1\leq i<j<k\leq 5$ we have that the lines $V_{i}$, $V_{j}$, $V_{k}$ are coplanar. Recalling that $V_{i}\cap U=\emptyset$, we have that the plane $\langle V_{i},V_{j},V_{k}\rangle$ meets $U$ at a point not contained in the union of $V_{i}$, $V_{j}$, and $V_{k}$. Examining Lemma 3.5, we have that $X_{f}(\mathbb{F}_{q})^{\mathrm{sm}}\neq\emptyset$.
Assume now that for no $(i,j,k)$ triples are $V_{i},V_{j},V_{k}$ coplanar, and set $u_{ij}=U\cap\langle V_{i},V_{j}\rangle$ for all $i<j$. The points $u_{12},u_{13},u_{23}$ are distinct, since otherwise $V_{1},V_{2},V_{3}$ are coplanar, so $f$ must not vanish on the line between one of them and $v$, by Lemma 3.6. Without loss of generality, assume $\langle u_{12},v\rangle\not\subset X_{f}$.
Consider the plane spanned by $V_{1},V_{2}$. Here $f|_{\langle V_{1},V_{2}\rangle}$ vanishes on the two lines $V_{1}$ and $V_{2}$, as well as a point $u_{12}$. By Lemma 3.5, either we have $X_{f}(\mathbb{F}_{q})^{\mathrm{sm}}\neq\emptyset$, or $f|_{\langle V_{1},V_{2}\rangle}$ vanishes on a line defined over $\overline{\mathbb{F}_{q}}$ containing $u_{12}$. In the latter case, if the line is not defined over $\mathbb{F}_{q}$, then in fact there exist two or three conjugate lines of multiplicity 1, intersecting at $u_{12}$. This allows us to choose distinct points $u_{i}\in V_{i}(\mathbb{F}_{q})-\{v\}$ for $i=1,2$ such that $u_{12}\notin\langle u_{1},u_{2}\rangle$. Since $\langle u_{1},u_{2}\rangle$ meets the aforementioned conjugate lines in distinct $\overline{\mathbb{F}_{q}}$-points, this satisfies the desired conditions.
Finally, suppose that $u_{12}$ is contained in a line $L\subset X_{f}\cap\langle V_{1},V_{2}\rangle$ defined over $\mathbb{F}_{q}$. We have $L\cap V_{i}=\emptyset$ for $i\neq 1,2$, since no three $V_{i}$ are coplanar and $v\notin L$. Thus for $i=3,4,5$, each $V_{i}$ meets the plane $\langle L,U\rangle$ at a single point $v_{i}$; moreover, these three points are not colinear, since the $V_{i}$ were not coplanar. Thus in the plane $\langle L,U\rangle$, $f|_{\langle L,U\rangle}$ vanishes on two lines and three non-colinear points, so by Lemma 3.5 we must have $X_{f}(\mathbb{F}_{q})^{\mathrm{sm}}\neq\emptyset$, and we are done.
Case $s\geq 2$. As in the previous case, by Lemma 4.4, there exist $v,w\in X_{f}(\mathbb{F}_{q})$ such that $\langle v,w\rangle\not\subset X_{f}$ and is skew to $U$. First, we argue there exists a chain of linear subvarieties defined over $\mathbb{F}_{q}$,
$u=U^{(s)}\subset U^{(s-1)}\subset\ldots\subset U^{(2)}\subset U^{(1)}\subset U^{(0)}=U,$
satisfying $\dim U^{(j)}=s-j$ and
(4.1) $\langle U^{(j)},x\rangle\not\subset X_{f}\text{ for all }x\in(X_{f}\cap\langle v,w\rangle)(\mathbb{F}_{q}).$
Here the superscript serves to remind us that $U^{(j)}\subset U$ has codimension $j$. To find $U^{(1)}$, we recognize that $\langle U,x\rangle\not\subset X_{f}$ for any $x\in(X_{f}-U)(\mathbb{F}_{q})$, since $\dim U$ is maximal, so applying Lemma 3.6 there are at most two possible codimension 1 $U^{(1)}\subset U$ defined over $\mathbb{F}_{q}$ for which $\langle U^{(1)},x\rangle\subset X_{f}$ for each $x\in(X_{f}\cap\langle v,w\rangle)(\mathbb{F}_{q})$. Since there are at most three $x\in(X_{f}\cap\langle u,v\rangle)(\mathbb{F}_{q})$ by Lemma 3.2, and more than 6 choices of $U^{(1)}$ by our hypotheses on $q$, there exists at least one satisfying (4.1). Repeating this argument, we find $U^{(j+1)}\subset U^{(j)}$ satisfying (4.1).
Suppose now that in $\langle U^{(1)},v,w\rangle\simeq\mathbb{P}^{s+1}$, $f$ vanishes on no linear subvarieties of dimension $s$. Then if $s>2$, the maximal dimension of such is s1s-1, and we may replace $U$ by $U^{(1)}$ and repeat the argument of this case. If $s=2$, then we have reduced to the case of $s=1$ above, so we are done.
Suppose to the contrary that for all of the $q+1$ codimension 1 $U_{i}^{(1)}\subset U$ defined over $\mathbb{F}_{q}$ satisfying $U^{(2)}\subset U_{i}^{(1)}$ (which implies (4.1) for $U_{i}^{(1)}$), we have that $f|_{\langle U_{i}^{(1)},v,w\rangle}$ vanishes on a linear subvariety $V_{i}$ of dimension $s$. Each $V_{i}$ meets the line $\langle v,w\rangle$ at exactly one of the at most three $\mathbb{F}_{q}$-points of $X_{f}\cap\langle v,w\rangle$. Thus whenever q>24q>{24}, there exist nine such $V_{i}$ containing the same $\mathbb{F}_{q}$-point of $X_{f}\cap\langle v,w\rangle$. Without loss of generality, assume $v\in V_{1},\ldots,V_{9}$.
Since $\dim(V_{i}\cap\langle u,v,w\rangle)=1$, we must have that $f$ vanishes on a line containing $v$ in $\langle u,v,w\rangle$. This line must meet $\langle u,w\rangle$, but by hypothesis $\langle u,v\rangle,\langle v,w\rangle\not\subset X_{f}$, so there must exist $y\in(X_{f}\cap\langle u,w\rangle)(\mathbb{F}_{q})$ such that $\langle v,y\rangle\subset V_{i}$ for all $i$. Note also that the line $\langle v,y\rangle\cap U=\emptyset$.
If $s>2$, let $V_{i}^{(3)}=V_{i}\cap U^{(2)}$, which has dimension s3s-3.333Here again the superscript serves to remind us that the codimension of $V_{i}^{(3)}\subset V_{i}$ is 3. Since $f$ vanishes on the span with $v$ of at most two codimension 1 linear subvarieties of $U^{(2)}$ by Lemma 3.6, at least five of the $V_{i}^{(3)}$ must coincide. Assume that for $V_{1},\ldots,V_{5}$ we have $V_{i}^{(3)}=V^{(3)}$. Now we have that $V^{(1)}=\langle V^{(3)},v,y\rangle\subset V_{i}$ for $i=1,\ldots,5$, so in fact the five $V_{i}$ meet in codimension 1. If $s=2$, then $V^{(3)}=\emptyset$ and $V^{(1)}=\langle v,y\rangle\subset V_{i}$ for all $i$, so we reach the same conclusion.
[…]
Our immediate goal is to use $V_{1}$, $V_{2}$, and $U$ to produce a plane $P$ which is skew to $V^{(1)}$ on which $f|_{P}\neq 0$ and vanishes on two lines. Then $P$ meets $V_{3},V_{4},V_{5}$ at three distinct points, and moreover these three points must not be colinear, lest $V_{3},V_{4},V_{5}$ span a space of dimension $s+1$, contradicting the discussion above. A final application of Lemma 3.5 then yields $X_{f}(\mathbb{F}_{q})^{\mathrm{sm}}\neq\emptyset$.
Let $v_{i}\in(V_{i}\cap U)-V^{(1)}$ for $i=1,2$. Thus we have $V_{i}\cap U=\langle V^{(1)}\cap U,v_{i}\rangle$ and $\langle v_{1},v_{2}\rangle\cap V^{(1)}=\emptyset$. Note that $\langle v_{1},v_{2}\rangle\subset X_{f}$, so applying Lemma 3.5 to $\langle V_{1},V_{2}\rangle$, we must have either that $f|_{\langle V_{1},V_{2}\rangle}$ vanishes on a smooth $\mathbb{F}_{q}$-point, vanishes on conjugate hyperplanes in $\langle V_{1},V_{2}\rangle$ that meet at a linear subvariety defined over $\mathbb{F}_{q}$ containing $\langle v_{1},v_{2}\rangle$, or vanishes on a third hyperplane $W\subset\langle V_{1},V_{2}\rangle$ containing $\langle v_{1},v_{2}\rangle$.
In the conjugate hyperplane case, we can find $\mathbb{F}_{q}$-points $u_{1}\in V_{1}-V^{(1)}$ and $u_{2}\in V_{2}-V^{(1)}$ not contained in the intersection of the conjugate hyperplanes. The line between them meets the conjugate hyperplanes at simple (geometric) points, so we are done.
Suppose now $W\subset\langle V_{1},V_{2}\rangle\cap X_{f}$ is defined over $\mathbb{F}_{q}$ with $\dim W=s$. Let $v_{2}^{\prime}\in(W-V^{(1)}-U)(\mathbb{F}_{q})$ so we have $\langle v_{1},v_{2}^{\prime}\rangle\subset X_{f}$ with the line meeting $U$ at $v_{1}$. By the maximality of $U$ and q>7q>{7}, successive applications of Lemma 3.6 show there exists $u^{\prime\prime}\in(U-\cup_{i=1}^{5}\langle V_{i},v_{1}\rangle)(\mathbb{F}_{q})$. Taking $P=\langle v_{1},v_{2}^{\prime},u^{\prime\prime}\rangle$, we see that $P\cap V^{(1)}=\emptyset$, while $f|_{P}$ vanishes on the lines $\langle v_{1},v_{2}^{\prime}\rangle\subset W$ and $\langle v_{1},u^{\prime\prime}\rangle\subset U$. By the discussion above, we have $X_{f}(F_{q})^{\mathrm{sm}}\neq\emptyset$, completing the proof. ∎
5. Proof of Theorem C
Proof of Theorem C.
Let $F\in\mathcal{O}_{K}[x_{0},\ldots,x_{n}]$ be a reduced form and denote by $f\in\mathbb{F}_{q}[x_{0},\ldots,x_{m}]$ its image under the residue map. If $X_{f}(\mathbb{F}_{q})^{\mathrm{sm}}\neq\emptyset$ then $X_{F}(K)\neq\emptyset$ by Hensel’s lemma.
Assume $X_{f}(\mathbb{F}_{q})^{\mathrm{sm}}=\emptyset$. By Proposition 4.5, we find $u,v\in X_{f}(\mathbb{F}_{q})$ such that $\langle u,v\rangle\not\subset X_{f}$ and $f|_{\langle u,v\rangle}$ vanishes on a point of multiplicity 1, necessarily defined over $\overline{\mathbb{F}_{q}}$ but not $\mathbb{F}_{q}$.
Possibly after a change of coordinates in $\operatorname{GL}_{n+1}(\mathcal{O}_{K})$, we may assume the $\overline{\mathbb{F}_{q}}$-points of $X_{f}$ along the line $\langle u,v\rangle$ are in the $x_{0}\neq 0$ affine patch. This allows us to find $(\mathbf{v}_{0},\mathbf{w}_{0})\in\mathbb{F}_{q}^{2n-1}$ such that
$f_{\mathbf{v}_{0},\mathbf{w}_{0},z}(X,0)=f(1,X+v_{1},w_{2}X+v_{2},\ldots,w_{n}X+v_{n})$
agrees with $f|_{\langle u,v\rangle}$. In particular, it vanishes on two distinct (singular) $\mathbb{F}_{q}$-points.
By Corollary 2.7, the hypothesis q>679q>{679} more than suffices to guarantee the existence of $\mathbf{z}_{0}$ such that $f_{\mathbf{v}_{0},\mathbf{w}_{0},\mathbf{z}_{0}}(X,Y)$ has no geometric linear factors through the marked (geometric) point along the $Y=0$ line.
Re-homogenizing, $f_{\mathbf{v}_{0},\mathbf{w}_{0},\mathbf{z}_{0}}(X,Y)$ corresponds to a degree 7 plane curve $C$ obtained by intersecting $X_{f}$ with a plane defined over $\mathbb{F}_{q}$. We have ensured that $C$ has a geometrically irreducible component defined over $\mathbb{F}_{q}$, namely that containing the $\overline{\mathbb{F}_{q}}$-point of multiplicity 1 on $\langle u,v\rangle$. Moreover, $C$ has at least two $\mathbb{F}_{q}$-points, $u$ and $v$.
If $C$ is geometrically reducible, its geometrically irreducible factor has degree $d^{\prime}$ at most 5.444If $C$ has a factor of degree 6 then it also has a linear factor over $\mathbb{F}_{q}$, so we find a smooth point. Letting $C^{\prime}$ denote this irreducible factor, with genus $g^{\prime}$, we apply Lemma 3.3, finding that q>159q>{159} suffices to ensure $C(\mathbb{F}_{q})^{\mathrm{sm}}\neq\emptyset$ (the limiting case is $C^{\prime}$ a smooth quintic curve).
It remains to consider the case of $C$ an irreducible degree 7 curve with at least two $\mathbb{F}_{q}$-points. If either of these points are smooth, we are done. If they are both singular, then the geometric genus is at most 13 (a smooth plane degree 7 plane curve has genus 15). Again applying (3.1), with $d^{\prime}=7$ and $g^{\prime}\leq 13$, we find $C(\mathbb{F}_{q})^{\mathrm{sm}}>0$ whenever q>679q>{679}.
Under the hypotheses q>679q>{679} and $F$ reduced, we have $X_{f}(\mathbb{F}_{q})^{\mathrm{sm}}\neq\emptyset$, therefore $X_{F}(K)\neq\emptyset$. It is enough to check this only on reduced forms $F$ by Proposition 4.3(i), so we conclude $K$ is $C_{2}(7)$ when q>679q>{679}. ∎